# AAPO: Enhance the Reasoning Capabilities of LLMs with Advantage Momentum

### v1

@@ Abstract
Reinforcement learning (RL) has emerged as an effective approach for enhancing the reasoning capabilities of large language models (LLMs), especially in scenarios where supervised fine-tuning (SFT) falls short due to limited chain-of-thought (CoT) data. Among RL-based post-training methods, group relative advantage estimation, as exemplified by Group Relative Policy Optimization (GRPO), has attracted considerable attention for eliminating the dependency on the value model, thereby simplifying training compared to traditional approaches like Proximal Policy Optimization (PPO). However, we observe that exsiting group relative advantage estimation method still suffers from training inefficiencies, particularly when the estimated advantage approaches zero. To address this limitation, we propose Advantage-Augmented Policy Optimization (AAPO), a novel RL algorithm that optimizes the cross-entropy (CE) loss using advantages enhanced through a momentum-based estimation scheme. This approach effectively mitigates the inefficiencies associated with group relative advantage estimation. Experimental results on multiple mathematical reasoning benchmarks demonstrate the superior performance of AAPO.
@@ 1 Introduction
Reinforcement learning (RL) has emerged as a powerful approach for enhancing the reasoning and decision-making capabilities of large language models (LLMs). While LLMs have demonstrated strong performance in both language understanding and generation tasks [3, 6, 40, 49], traditional training strategies such as pre-training and supervised fine-tuning (SFT) [30, 2, 24] often fall short in enabling effective chain-of-thought (CoT) [43] reasoning for complex decision-making tasks. To address this limitation, recent research has explored RL-based training paradigms, which have shown considerable empirical success in specialized domains such as mathematical reasoning. Models including GPT-o1 [26], DeepSeek-R1 [14], and QwQ [29] exemplify this promising direction, demonstrating the potential of RL to substantially improve the reasoning abilities of LLMs.
@@
A recent advancement in LLM post-training with RL is the introduction of novel methods for advantage estimation, a technique for quantifying how favorable a specific action is in a given state. In this context, the term advantage measures the relative benefit of an action compared to the average in that state, thereby providing an informative learning signal during training. Traditionally, approaches such as Proximal Policy Optimization (PPO) [35], exemplified by InstructGPT [27], rely on a value model to estimate the advantage. Although PPO offers stable and reliable performance, the need to maintain a separate value model leads to substantial consumption of GPU resources.
@@
In contrast to conventional approaches, the group relative advantage estimation method was originally proposed in Group Relative Policy Optimization (GRPO) [36] and has since been widely adopted for enhancing reasoning capabilities in LLMs. Group relative advantage estimation method removes the need for value models entirely by evaluating responses relative to the average within a group of sampled responses. This approach significantly reduces GPU memory usage and computational costs while maintaining competitive performance in downstream reasoning tasks. Its effectiveness is further evidenced by the strong performance of modern reasoning models such as DeepSeek-R1 [14], which highlights the effectiveness of group relative advantage estimation in balancing computational efficiency and performance robustness. Several extensions have further refined this paradigm. Decoupled Clipping and Dynamic sAmpling Policy Optimization (DAPO) [48] improves training on long CoT sequences through token-level gradient estimation and relaxed clipping strategies. Dr. GRPO [25] introduces an unbiased optimization method that enhances token efficiency, while Group Policy Gradient (GPG) [8] further simplifies the learning process by removing surrogate losses and eliminating the reference model. While these methods vary in implementation specifics, they all share a common foundation in the principle of group relative advantage estimation for effective LLMs post-training. We provide a comprehensive discussion of GRPO and GPG in Section 3.
@@
Although the above approaches have significantly advanced RL in the post-training stage and enhanced the reasoning capabilities of LLMs beyond what SFT can achieve, the practical limitations associated with the group relative advantage estimation method remain unresolved. As shown in the left panel of Figure 1, when the group relative advantage estimation method [36, 8, 25] is adopted, the advantage may approach when the rewards within a group exhibit low variance, resulting in zero gradient and, consequently, no parameter updates. Conversely, when rewards among responses within the group vary significantly, the resulting advantages can exhibit high variance, potentially leading to unstable or unintended gradient ascent. Both scenarios deviate from the desired optimization trajectories.
@@
Figure 1: Left: An example convergence trajectories of policy optimization with group relative advantage estimation method proposed in GRPO. Red node denotes gradient euqals to zero. Purple arrow denotes gradient ascent. Green arrow denotes gradient descent. Right: An example of ideal convergence trajectories of policy optimization.
@@
  • We propose Advantage-Augmented Policy Optimization (AAPO), which mitigates the issues of advantage estimation with advantage momentum by taking a comparison with the reference model.
  • Numerous experiments have demonstrated that AAPO exhibits better optimization behavior during training and achieves superior performance on representative mathematical reasoning benchmarks.
@@ 2 Related Work
Large Language Models LLMs such as PaLM 2 [1] and Llama [40, 41] achieved impressive performance in understanding and generation tasks, largely due to scaling laws, high-quality pre-train corpus, and optimized training techniques [5, 27, 13, 42]. However, these models still struggle with complex reasoning. In contrast, recent LLMs such as GPT-o1 [26], DeepSeek-R1 [14], and QwQ [29] show strong reasoning capabilities, particularly in mathematical tasks. These gains are primarily attributed to RL-based post-training, which significantly outperforms traditional SFT.
Reinforcement Learning for LLMs  Early RL-based post-training methods [7, 37, 27, 50] mainly relied on human-labeled preference data and reward models to evaluate responses. PPO [35] estimates state values via a value model for advantage estimation while using a reward model to assign rewards. DPO [31] uses paired preference data to encourage preferred responses and suppress undesired ones. However, annotating preferences and training reward models are resource intensive. Moreover, the scarcity of explicit reasoning data limits the enhancement of models’ reasoning capabilities via these RL methods. DeepSeek-R1 is the first to report the use of RL to enable extended CoT generation and the emergence of the so-called "aha moment" [14]. Specifically, DeepSeek-R1 adopts GRPO [36], which estimates advantages through in-group comparisons, thereby enhancing the feasibility of aligning models for reasoning tasks. Yet, the limitations of existing advantage estimation remain underexplored. In this work, we introduce AAPO, which redefines advantage estimation by incorporating advantage momentum, effectively addressing the issues of the group relative advantage estimation method in policy optimization.
Advantage estimation in Reinforcement Learning  Previous RL approaches [35, 33, 15] estimate the advantage using either Monte Carlo returns [45], Temporal-Difference (TD) [38] errors, or Generalized Advantage Estimation (GAE) [34]. While these critic-based estimators form the backbone of algorithms such as PPO, maintaining a separate value network becomes computationally expensive when the policy is implemented as an LLM. Group relative advantage estimation method proposed by GRPO [36] computes the advantage by comparing each response’s reward to the group mean, reducing the resources needed. However, it introduces new optimization challenges, as discussed in the introduction section, that have yet to be investigated. In this work, we propose a novel advantage estimation method that incorporates advantage momentum and optimizes the cross-entropy loss to better enhance the reasoning capabilities of LLMs.
@@ 3 Preliminary
This section provides a preliminary overview of group-based relative advantage estimation methods for RL-based training by briefly introducing GRPO [36] and GPG [8]. In the context of the GRPO algorithm, it circumvents the dependency on the value model which is commonly required in PPO [35] by estimating advantage within grouped samples. It utilizes a rule-based reward system to score responses generated by LLMs. Furthermore, GRPO retains the clipping strategy from PPO to prevent excessively large policy updates and leverages the reference model $\pi_{ref}$ to compute the current Kullback-Leibler (KL) divergence, thereby ensuring the stability and integrity of the model during training. Specifically, for each question $q$, GRPO samples a group of responses $O=\{o_{1},o_{2},\cdots,o_{G}\}$ from the old policy $\pi_{\theta_{\mathrm{old}}}$ and optimizes the policy model $\pi_{\theta}$ by maximizing the following objective:
@@
$\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t}^{GRPO},\operatorname{clip}\left(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}^{GRPO}\right]-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},$
where $\varepsilon$ and $\beta$ are hyper-parameters to control the clip boundary and the KL divergence penalty coefficient, respectively. In this context, $\hat{A}_{i,t}^{GRPO}$ represents the advantage, derived exclusively from the relative reward of the responses within the same group. $r_{i,t}(\theta)$ represents the likelihood ratio between the current policy $\pi_{\theta}$ and the old policy $\pi_{\theta_{\mathrm{old}}}$. Typically, the likelihood ratio $r_{i,t}(\theta)$ and the advantage $\hat{A}_{i,t}^{GRPO}$ are calculated using the following formula:
$r_{i,t}(\theta)=\frac{\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{i,t}\mid q,o_{i,<t}\right)},\quad\hat{A}_{i,t}^{GRPO}=\frac{r_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}.$ (2)
Recent work GPG [8] proposes directly optimizing the original RL objective, yielding improved performance over GRPO. While the advantage estimation method used in GPG is similar to that of GRPO, it further emphasizes the advantage of responses that are valid to gradient estimation. Generally speaking, the core objective of GPG is to optimize the following objective:
$\mathcal{J}_{\mathrm{GPG}}(\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D},\left\{o_{i}\right\}_{i=1}^{G}}{\left[\frac{1}{\sum_{i=1}^{G}\left|o_{i}\right|}\sum_{i=1}^{G}\sum_{t=1}^{\left|o_{i}\right|}\left(-\log\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)\hat{A}_{i,t}\right)\right],}$ (3)
where $\hat{A}_{i,t}^{GPG}=\frac{r_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{F_{norm}}$ and $F_{norm}$ could be 1 or $\operatorname{std}(\{R_{i}\}_{i=1}^{G})$. However, the rewards of the responses within a group may exhibit low variance. This suggests that $\hat{A}_{i,t}^{GPG}$ in GPG encounters the same challenges as $\hat{A}_{i,t}^{GRPO}$ in GRPO.
In this work, we propose a novel algorithm AAPO to address the issues caused by the group relative advantage estimation method in the policy optimization process. We also provide an in-depth analysis of both the prevailing advantage estimation method adopted in [36, 48, 8] and our proposed AAPO.
@@ 4 Advantage-Augmented Policy Optimization
Inspired by the optimization behavior found in current RL algorithms [36, 25, 48, 8], our objective is to mitigate the issues that arise in the policy optimization process, overcoming the challenge where advantage estimation tends to approach zero or bad advantage estimation in the later steps of RL training. Drawing further inspiration from the well-established Adam [21] and the recent GPG algorithm [8], which optimizes the RL objective directly, thus avoiding the surrogate loss function, we propose a novel algorithm, Advantage-Augmented Policy Optimization (AAPO), which directly optimizes the cross-entropy (CE) loss enhanced by augmented advantage, which is driven by the advantage momentum. In contrast to previous approaches [36, 8, 25], AAPO leverages advantage amplification by performing group-based sampling for both the policy model $\pi_{\theta}$ and the reference model $\pi_{ref}$. Specifically, we calculate the reward for each response generated by the policy model $\pi_{\theta}$, evaluate the relative advantage of each response within its group $G$, and compare these rewards $r_{\theta_{i}}$ with those $r_{ref_{i}}$ obtained from the reference model $\pi_{ref}$. This method improves the effectiveness of the RL training step by preventing the advantage from approaching zero and exhibiting high variance. Formally, AAPO optimizes the policy model $\pi_{\theta}$ by minimizing the following objective:
@@
Based on the optimization objective in Equation (4), it is straightforward to derive its gradient:

### v2

# AAPO: Enhancing the Reasoning Capabilities of LLMs with Advantage Momentum

@@ Abstract
Reinforcement learning (RL) has emerged as an effective approach for enhancing the reasoning capabilities of large language models (LLMs), especially in scenarios where supervised fine-tuning (SFT) falls short due to limited chain-of-thought (CoT) data. Among RL-based post-training methods, group relative advantage estimation, as exemplified by Group Relative Policy Optimization (GRPO), has attracted considerable attention for eliminating the dependency on the value model, thereby simplifying training compared to traditional approaches like Proximal Policy Optimization (PPO). However, we observe that existing group relative advantage estimation method still suffers from training inefficiencies, particularly when the estimated advantage approaches zero. To address this limitation, we propose Advantage-Augmented Policy Optimization (AAPO), a novel RL algorithm that optimizes the cross-entropy (CE) loss using advantages enhanced through a momentum-based estimation scheme. This approach effectively mitigates the inefficiencies associated with group relative advantage estimation. Experimental results on multiple mathematical reasoning benchmarks and model series demonstrate the superior performance of AAPO.
@@ 1 Introduction
Reinforcement learning (RL) has emerged as a powerful approach for enhancing the reasoning and decision-making capabilities of large language models (LLMs). While LLMs have demonstrated strong performance in both language understanding and generation tasks (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023a; Zhao et al., 2023), traditional training strategies such as pre-training and supervised fine-tuning (SFT) (Radford et al., 2018; Bommasani et al., 2021; Liu et al., 2023) often fall short in enabling effective chain-of-thought (CoT) (Wei et al., 2022) reasoning for complex decision-making tasks. To address this limitation, recent research has explored RL-based training paradigms, which have shown considerable empirical success in specialized domains such as mathematical reasoning. Models including GPT-o1 (OpenAI, 2024), DeepSeek-R1 (Guo et al., 2025), and QwQ (Qwen Team, 2024) exemplify this promising direction, demonstrating the potential of RL to substantially improve the reasoning abilities of LLMs.
@@
A recent advancement in LLM post-training with RL is the introduction of novel methods for advantage estimation, a technique for quantifying how favorable a specific action is in a given state. In this context, the term advantage measures the relative benefit of an action compared to the average in that state, thereby providing an informative learning signal during training. Traditionally, approaches such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), exemplified by InstructGPT (Ouyang et al., 2022), rely on a value model to estimate the advantage. Although PPO offers stable and reliable performance, the need to maintain a separate value model leads to substantial consumption of GPU resources.
@@
In contrast to conventional approaches, the group relative advantage estimation method was originally proposed in Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and has since been widely adopted for enhancing reasoning capabilities in LLMs. Group relative advantage estimation method removes the need for value models entirely by evaluating responses relative to the average within a group of sampled responses. This approach significantly reduces GPU memory usage and computational costs while maintaining competitive performance in downstream reasoning tasks. Its effectiveness is further evidenced by the strong performance of modern reasoning models such as DeepSeek-R1 (Guo et al., 2025), which highlights the effectiveness of group relative advantage estimation in balancing computational efficiency and performance robustness. Several extensions have further refined this paradigm. Decoupled Clipping and Dynamic sAmpling Policy Optimization (DAPO) (Yu et al., 2025) improves training on long CoT sequences through token-level gradient estimation and relaxed clipping strategies. Dr. GRPO (Liu et al., 2025) introduces an unbiased optimization method that enhances token efficiency, while Group Policy Gradient (GPG) (Chu et al., 2025) further simplifies the learning process by removing surrogate losses and eliminating the reference model. While these methods vary in implementation specifics, they all share a common foundation in the principle of group relative advantage estimation for effective LLMs post-training. We provide a comprehensive discussion of GRPO and GPG in Section 3.
@@
Although the above approaches have significantly advanced RL in the post-training stage and enhanced the reasoning capabilities of LLMs beyond what SFT can achieve, the practical limitations associated with the group relative advantage estimation method remain unresolved. As shown in the left panel of Figure 1, when the group relative advantage estimation method (Shao et al., 2024; Chu et al., 2025; Liu et al., 2025) is adopted, the advantage may approach when the rewards within a group exhibit low variance, resulting in zero gradient and, consequently, no parameter updates. Conversely, when rewards among responses within the group vary significantly, the resulting advantages can exhibit high variance, potentially leading to unstable or unintended gradient ascent. Both scenarios deviate from the desired optimization trajectories.
@@
Figure 1: Left: An example convergence trajectories of policy optimization with group relative advantage estimation method proposed in GRPO. Red node denotes gradient equals to zero. Purple arrow denotes gradient ascent. Green arrow denotes gradient descent. Right: An example of ideal convergence trajectories of policy optimization.
@@
  • We propose Advantage-Augmented Policy Optimization, which mitigates the issues of advantage estimation with advantage momentum by taking a comparison with the reference model.
  • Experiments have demonstrated that AAPO achieves superior performance across different model series and mathematical reasoning benchmarks.
@@ 2 Related work
Large Language Models  LLMs such as PaLM 2 (Anil et al., 2023) and Llama (Touvron et al., 2023a; b) achieved impressive performance in understanding and generation tasks, largely due to scaling laws, high-quality pre-train corpus, and optimized training techniques (Chen et al., 2024; Ouyang et al., 2022; Gunel et al., 2021; Wang et al., 2023). However, these models still struggle with complex reasoning. In contrast, recent LLMs such as GPT-o1 (OpenAI, 2024), DeepSeek-R1 (Guo et al., 2025), and QwQ (Qwen Team, 2024) show strong reasoning capabilities, particularly in mathematical tasks. These gains are primarily attributed to RL-based post-training, which significantly outperforms traditional SFT.
Reinforcement Learning for LLMs  Early RL-based post-training methods (Christiano et al., 2017; Song et al., 2024; Ouyang et al., 2022; Ziegler et al., 2019) mainly relied on human-labeled preference data and reward models to evaluate responses. PPO (Schulman et al., 2017) estimates state values via a value model for advantage estimation while using a reward model to assign rewards. DPO (Rafailov et al., 2023) uses paired preference data to encourage preferred responses and suppress undesired ones. However, annotating preferences and training reward models are resource intensive. Moreover, the scarcity of explicit reasoning data limits the enhancement of models’ reasoning capabilities via these RL methods. DeepSeek-R1 is the first to report the use of RL to enable extended CoT generation and the emergence of the so-called ”aha moment” (Guo et al., 2025). Specifically, DeepSeek-R1 adopts GRPO (Shao et al., 2024), which estimates advantages through in-group comparisons, thereby enhancing the feasibility of aligning models for reasoning tasks. Yet, the limitations of existing advantage estimation remain underexplored. In this work, we introduce AAPO, which redefines advantage estimation by incorporating advantage momentum, effectively addressing the issues of the group relative advantage estimation method in policy optimization.
Advantage estimation in Reinforcement Learning  Previous RL approaches (Schulman et al., 2017; 2015; Haarnoja et al., 2018) estimate the advantage using either Monte Carlo returns (Williams, 1992), Temporal-Difference (TD) (Sutton et al., 1999) errors, or Generalized Advantage Estimation (GAE) (Schulman et al., 2016). While these critic-based estimators form the backbone of algorithms such as PPO, maintaining a separate value network becomes computationally expensive when the policy is implemented as an LLM. Group relative advantage estimation method proposed by GRPO (Shao et al., 2024) computes the advantage by comparing each response’s reward to the group mean, reducing the resources needed. However, it introduces new optimization challenges, as discussed in the introduction section, that have yet to be investigated. In this work, we propose a novel advantage estimation method that incorporates advantage momentum and optimizes the cross-entropy loss to better enhance the reasoning capabilities of LLMs.
@@ 3 Preliminary
This section provides a preliminary overview of group-based relative advantage estimation methods for RL-based training by briefly introducing GRPO (Shao et al., 2024) and GPG (Chu et al., 2025). In the context of the GRPO algorithm, it circumvents the dependency on the value model which is commonly required in PPO (Schulman et al., 2017) by estimating advantage within grouped samples. It utilizes a rule-based reward system to score responses generated by LLMs. Furthermore, GRPO retains the clipping strategy from PPO to prevent excessively large policy updates and leverages the reference model $\pi_{ref}$ to compute the current Kullback-Leibler (KL) divergence, thereby ensuring the stability and integrity of the model during training. Specifically, for each question $q$, GRPO samples a group of responses $O=\{o_{1},o_{2},\cdots,o_{G}\}$ from the old policy $\pi_{\theta_{\mathrm{old}}}$ and optimizes the policy model $\pi_{\theta}$ by maximizing the following objective:
@@
$\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t}^{\mathrm{GRPO}},\operatorname{clip}\left(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}^{\mathrm{GRPO}}\right]-\beta\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},$
where $\varepsilon$ and $\beta$ are hyper-parameters to control the clip boundary and the KL divergence penalty coefficient, respectively. In this context, $\hat{A}_{i,t}^{\mathrm{GRPO}}$ represents the advantage, derived exclusively from the relative reward of the responses within the same group. $r_{i,t}(\theta)$ represents the likelihood ratio between the current policy $\pi_{\theta}$ and the old policy $\pi_{\theta_{\mathrm{old}}}$. Typically, the likelihood ratio $r_{i,t}(\theta)$ and the advantage $\hat{A}_{i,t}^{\mathrm{GRPO}}$ are calculated using the following formula:
$r_{i,t}(\theta)=\frac{\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{i,t}\mid q,o_{i,<t}\right)},\quad\hat{A}_{i,t}^{\mathrm{GRPO}}=\frac{r_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}.$ (2)
Recent work GPG (Chu et al., 2025) proposes directly optimizing the original RL objective, yielding improved performance over GRPO. While the advantage estimation method used in GPG is similar to that of GRPO, it further emphasizes the advantage of responses that are valid to gradient estimation. Generally speaking, the core objective of GPG is to optimize the following objective:
$\mathcal{J}_{\mathrm{GPG}}(\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D},\left\{o_{i}\right\}_{i=1}^{G}}{\left[\frac{1}{\sum_{i=1}^{G}\left|o_{i}\right|}\sum_{i=1}^{G}\sum_{t=1}^{\left|o_{i}\right|}\left(-\log\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)\hat{A}_{i,t}^{\mathrm{GPG}}\right)\right],}$ (3)
where $\hat{A}_{i,t}^{\mathrm{GPG}}=\frac{r_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{F_{norm}}$ and $F_{norm}$ could be 1 or $\operatorname{std}(\{R_{i}\}_{i=1}^{G})$. However, the rewards of the responses within a group may exhibit low variance. This suggests that $\hat{A}_{i,t}^{\mathrm{GPG}}$ in GPG encounters the same challenges as $\hat{A}_{i,t}^{\mathrm{GRPO}}$ in GRPO.
In this work, we propose a novel algorithm AAPO to address the issues caused by the group relative advantage estimation method in the policy optimization process. We also provide an in-depth analysis of both the prevailing advantage estimation method adopted in (Shao et al., 2024; Yu et al., 2025; Chu et al., 2025) and our proposed AAPO.
@@ 4 Advantage-Augmented Policy Optimization
Inspired by the optimization behavior found in current RL algorithms (Shao et al., 2024; Liu et al., 2025; Yu et al., 2025; Chu et al., 2025), our objective is to mitigate the issues that arise in the policy optimization process, overcoming the challenge where advantage estimation tends to approach zero or bad advantage estimation in the later steps of RL training. Drawing further inspiration from the well-established Adam (Kingma & Ba, 2015) and the recent GPG algorithm (Chu et al., 2025), which optimizes the RL objective directly, thus avoiding the surrogate loss function, we propose a novel algorithm, Advantage-Augmented Policy Optimization (AAPO), which directly optimizes the cross-entropy (CE) loss enhanced by augmented advantage, which is driven by the advantage momentum. In contrast to previous approaches (Shao et al., 2024; Chu et al., 2025; Liu et al., 2025), AAPO leverages advantage amplification by performing group-based sampling for both the policy model $\pi_{\theta}$ and the reference model $\pi_{ref}$. Specifically, we calculate the reward for each response generated by the policy model $\pi_{\theta}$, evaluate the relative advantage of each response within its group $G$, and compare these rewards $r_{\theta_{i}}$ with those $r_{ref_{i}}$ obtained from the reference model $\pi_{ref}$. This method improves the effectiveness of the RL training step by preventing the advantage from approaching zero and exhibiting high variance. Formally, AAPO optimizes the policy model $\pi_{\theta}$ by minimizing the following objective:
@@
Based on equation 4, it is straightforward to derive its gradient:
@@
DefinitionFor a group $\mathcal{G}$ containing sampled responses $O=\{o_{1},o_{2},\cdots,o_{\mathcal{G}}\}$, the empirical AAPO loss is defined as $\mathcal{L}_{\mathcal{G}}(\theta)=\frac{1}{N_{\mathcal{G}}}\sum_{o\in\mathcal{G}}[-\log\pi_{\theta}(o)\hat{A}^{*}]$, where $\pi_{\theta}$ is the policy model, $N_{\mathcal{G}}=\sum_{o\in\mathcal{G}}|o|$ is the total number of tokens in the group. We further define the expected objective as the expectation of the empirical loss over all possible groups $\mathcal{L}(\theta)=\mathbb{E}_{\mathcal{G}\sim\pi_{\theta}}[\mathcal{L}_{\mathcal{G}}(\theta)]$.
Theorem 1.
(Stability)  Since the rewards are bounded, the group standard deviation satisfies $0\leq\sigma_{min}\leq\sigma$, and the log-likelihood gradients are bounded as $\nabla_{\theta}\log\pi_{\theta}(o)||\leq M$. Then, Each gradient step with learning rate $\eta_{k}$ satisfies $\theta_{k+1}-\theta_{k}||\leq\eta_{k}MB$, where $B=\frac{R_{max}-R_{min}}{\sigma_{min}}+\operatorname{max}(|\delta_{\mathrm{low}}|,|\delta_{\mathrm{high}}|)$ is the uniform bound on the AAPO weights. The expected objective is bounded from $\mathcal{L}(\theta)\geq-B\log|\mathcal{V}|$, where $|\mathcal{V}|$ is the vocabulary size. Hence, AAPO training is stable: the objective cannot diverge to -$\infty$ and parameter updates are always finite. Proof in Appendix B.
Theorem 2.
(Convergence)  Assume that the stochastic gradient is unbiased and that the per-sample gradient has bounded second moment. Let the step sizes satisfy the Robbins–Monro conditions $\eta_{k}>0$, $\sum_{k}\eta_{k}=\infty$, $\sum_{k}\eta_{k}^{2}<\infty$. AAPO converges to a stationary point of its expected objective $\liminf_{k\to\infty}\mathbb{E}\left[\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]=0$. Moreover, if a constant step size $\eta<\frac{1}{BL_{0}}$ is used, where $L_{0}$ is the smoothness constant of $\log\pi_{\theta}(o)$, then the iterates converge to a neighborhood of stationarity $\limsup_{K\to\infty}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\!\left[\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]\;\lesssim\;\mathcal{O}(\eta)+\mathcal{O}\!\left(\tfrac{1}{N_{\mathcal{G}}}\right)$. Proof in Appendix B.